OpenStudy (anonymous):

Operations with Radical Expressions Simplify 7√7 - 2√7 Please explain all the steps, I have these for homework and I'm trying to understand how to solve them, thanks. =)

OpenStudy (asnaseer):

how would you simplify: 7x - 2x

OpenStudy (asnaseer):

the problem you stated is very similar to this

OpenStudy (asnaseer):

@LemonLeaf - do you understand?

OpenStudy (jdoe0001):

$$\large 2\sqrt{x}+3\sqrt{x} \ \color{red}{\text{ if a=}\sqrt{x}}\implies 2a+3a = \boxed{\ ?} $$

OpenStudy (anonymous):

so It would be 5√7?

OpenStudy (asnaseer):

perfect! :)

OpenStudy (asnaseer):

but do you understand why?

OpenStudy (anonymous):

not really

OpenStudy (asnaseer):

ok - let me try and explain...

OpenStudy (anonymous):

I suppose it's like simplifying with variables, but the variable being replaced with a square root

OpenStudy (asnaseer):

think of \(\sqrt{7}\) as some number. then \(7\sqrt{7}\) just means 7 times this number. so you can just replace \(\sqrt{7}\) by some symbol to represent its value - lets use 'x'. therefore \(7\sqrt{7}=7x\)

OpenStudy (asnaseer):

and yes, you statement is true - you can think of it in that way as well

OpenStudy (jdoe0001):

once the root element, that is the root and its content RESEMBLE the same, you can treat them as any variable and just work with their coefficients

OpenStudy (anonymous):

oh okay that's simple enough, thanks for your help :)

OpenStudy (asnaseer):

but you must be careful to replace like-with-like e.g. you couldn't simplify: \(7\sqrt{7}-2\sqrt{5}\) because \(\sqrt{7}\) and \(\sqrt{5}\) are two DIFFERENT quantities.

OpenStudy (asnaseer):

yw :)

OpenStudy (anonymous):

could you help me with another one? it's a bit different than this one

OpenStudy (asnaseer):

sure - just post it as a new question on the left please (after closing this one)

OpenStudy (jdoe0001):

if it were say \(\sqrt{2}\ and \ \sqrt{3} \) then that wouldn't work, but if you had \(\large \sqrt[5]{2} \ and \ \sqrt[5]{2}\), then they RESEMBLE each other, then you can